### v2

# Learning the Basis: A Kolmogorov-Arnold Network Approach Embedding Green’s Function Priors

Abstract
The Method of Moments (MoM) is constrained by the usage of static, geometry-defined basis functions, such as the Rao-Wilton-Glisson (RWG) basis. This letter reframes electromagnetic modeling around a learnable basis representation rather than solving for the coefficients over a fixed basis. We first show that the RWG basis is essentially a static and piecewise-linear realization of the Kolmogorov-Arnold representation theorem. Inspired by this insight, we propose PhyKAN, a physics-informed Kolmogorov-Arnold Network (KAN) that generalizes RWG into a learnable and adaptive basis family. Derived from the EFIE, PhyKAN integrates a local KAN branch with a global branch embedded with Green’s function priors to preserve physical consistency. It is demonstrated that, across canonical geometries, PhyKAN achieves sub-$10^{-2}$ reconstruction errors as well as accurate, unsupervised radar cross section predictions, offering an interpretable, physics-consistent bridge between classical solvers and modern neural network models for electromagnetic modeling.
I Introduction
Solving electromagnetic (EM) scattering via the Electric Field Integral Equation (EFIE) is central to computational electromagnetics (CEM) [1], with the Method of Moments (MoM) being the standard approach. This method expands the induced surface current using predefined basis functions, typically the Rao–Wilton–Glisson (RWG) basis [2]. However, these basis functions are fixed by mesh geometry and are inherently static, a fact that prevents them from adapting to the spatial complexity of current distributions. This rigidity reveals a key limitation of MoM, as its basis functions are not learnable and cannot adapt to the underlying physics.
To address the latter limitation, several directions have been explored. Higher-order basis functions enrich MoM expressiveness using higher-degree polynomials [3, 4, 5], while characteristic or macro-basis constructions improve efficiency through physics-informed domain decomposition [6, 7]. On another front, instead of improving the basis itself, Adaptive Mesh Refinement (AMR) refines the discretization in regions of high error [8, 9]. Although effective, these approaches remain largely hand-crafted, problem-specific, and fundamentally tied to predefined basis formulations.
Deep learning techniques are recently emerging as efficient tools in CEM problems, mainly along two directions: black-box current or field prediction without any explicit basis representation [10, 11], and MoM acceleration through surrogate models, operator learning, or iterative refinement, still relying on fixed basis functions [12, 13, 14]. More recent studies integrate physical structure into neural models, such as EFIE-guided graph learning [15, 16]. This indicates that neither direction addresses the non-learnable nature of the basis functions, which remain unable to adapt to EM physics.
Recent progress in scientific machine learning enables neural network architectures that integrate physical structure withing learnable representations. The Kolmogorov–Arnold Network (KAN) [17] is one of the well suited learning frameworks for this purpose that replaces fixed activations with learnable univariate functions [18], providing adaptable basis behavior. KAN is flexible in both kernel choices [19, 20] and graph-based extensions [21] that embed geometric and physical interactions, making it a strong foundation for a physics-guided learnable basis for the MoM approach.
In this letter, we reformulate MoM basis construction as a learnable function problem and show that KAN provides a theoretically grounded generalization of the RWG basis. In addition, a physics-guided EFIE-based framework, termed as PhyKAN, that embeds Green’s function [22] priors for physically consistent and adaptive current modeling, while remaining compact and interpretable, is introduced. The proposed learning framework accommodates multiple kernel functions and structural extensions, offering a unified bridge between classical integral equation solvers and neural network models.
II Surface Current as a Learning Problem
Figure 1: (a) The architecture of the proposed PhyKAN comprising a local branch that encodes geometric features using a conventional KAN and global branch that incorporates Green’s function priors for physically consistent prediction of the induced surface current. (b) The structure of the conventional KAN [17], where learnable activation functions enable adaptive basis modeling.
Let $\mathbf{G}\left(\cdot\right)\in\mathbb{C}^{3\times 3}$ denote the dyadic Green’s function in the three-dimensional free space, $\boldsymbol{J}(\cdot)\in\mathbb{C}^{3\times 1}$ the induced surface current, and $\boldsymbol{E}_{\rm inc}(\cdot)\in\mathbb{C}^{3\times 1}$ the incident field. By applying the boundary conditions of Perfectly Electrically Conducting (PEC) objects $\mathcal{D}$, the EFIE can be formulated $\forall\boldsymbol{p}\in\mathbb{R}^{3}$ as:
$\int_{\boldsymbol{p}^{\prime}\in\mathcal{D}}\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)\boldsymbol{J}\left(\boldsymbol{p}^{\prime}\right)d\boldsymbol{p}^{\prime}=\boldsymbol{E}_{\rm inc}(\boldsymbol{p}).$ (1)
The MoM discretizes this integral equation by expanding the unknown current $\boldsymbol{J}(\boldsymbol{p})$ using a set of predefined basis functions. Typically, the RWG basis $\{{f}_{n}(\cdot)\}_{n=1}^{3}$ comprising fixed and geometry-defined functions [2] is used, yielding the following surface current approximation:
$\boldsymbol{J}\left(\boldsymbol{p}_{k}^{\prime}\right)\approx\sum_{n=1}^{3}I_{n,k}{f}_{n}\left(\boldsymbol{p}_{k}^{\prime}\right),$ (2)
where $I_{n,k}$ are the unknown complex coefficients of the $n$-th RWG basis function on the $k$-th triangular element centered at $\boldsymbol{p}_{k}^{\prime}$. This geometry-fixed basis restricts the MoM’s adaptability to complex current distributions. Existing refinements, such as higher-order bases [3, 4, 5], the characteristic basis function method (CBFM) [7], and AMR[8, 9], partially alleviate this limitation but do not alter the static nature of the basis itself.
Although recent deep learning works [10, 11, 12, 13, 14, 15, 16] attempt to either learn geometry-to-current mappings or accelerate MoM via surrogate models, they still operate on fixed basis functions, thus leaving the aforedescribed core limitation unresolved. This reveals a fundamental gap: a formulation that preserves the MoM structure while enabling adaptive basis functions is still missing. To this end, instead of solving for $I_{n,k}$ in (2) satisfying the EFIE in (1), we focus on learning an optimal, data-driven basis family $\left\{b_{n}(\cdot;\boldsymbol{u})\right\}$, yielding the following generalized expansion:
$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum^{N_{b}}_{n=1}\alpha_{n,k}b_{n}(\boldsymbol{p}^{\prime}_{k};\boldsymbol{u}).$ (3)
Here, $N_{b}$ denotes the number of learnable basis functions used per element, $\alpha_{n,k}$ is the coefficient associated with the $n$-th learnable basis on the $k$-th triangular element, and $b_{n}(\cdot;\boldsymbol{u})$ is a learnable function parameterized by $\boldsymbol{u}$, whose goal is to adapt its functional shape locally based on the underlying physics and geometry. Notably, setting $b_{n}=f_{n}$ and $\alpha_{n,k}=I_{n,k}$ recovers the classical MoM expansion in (2) as a special case.
The latter problem finds its direct mathematical foundation in the Kolmogorov–Arnold representation theorem, which states that any multivariate function can be decomposed into finite sums of univariate functions [23]:
$f\left(x_{1},\ldots,x_{d}\right)=\sum_{i=1}^{2d+1}\alpha_{i}\left(\sum_{j=1}^{d}\phi_{ij}\left(x_{j}\right)\right),$ (4)
where $d$ is the input dimension, $\alpha_{i}$ are learnable weights, and $\phi_{ij}(\cdot)$’s are univariate nonlinear mappings. The KAN is the neural network architecture that operationalizes this theorem. To construct learnable basis functions, we embed the KAN parameterization into each $b_{n}(\cdot;\boldsymbol{u})$ as
$b_{n}\left(\boldsymbol{p}^{\prime}_{k};\boldsymbol{u}_{n}\right)=\sum_{i=1}^{M}c_{n,i}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right),$ (5)
where $\phi(\cdot)$ plays the same role as the univariate nonlinear mappings $\phi_{ij}(\cdot)$ in (4), and $\left\{\boldsymbol{w}_{n,i},\beta_{n,i},c_{n,i}\right\}$ are learnable parameters controlling each kernel’s orientation, offset, and amplitude, respectively. The term $\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right)$ performs a learnable linear projection of the spatial coordinate $\boldsymbol{p}^{\prime}_{k}$, allowing each kernel to adapt to different geometric directions and field variations.
Substituting (5) into (3), yields the approximation:
$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum_{n=1}^{N_{b}}\sum_{i=1}^{M}\alpha_{n,k}c_{n,i}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right).$ (6)
To more closely mirror the MoM form in (2), we next absorb the coefficient into the spline weights via the definition $\tilde{c}_{n,i,k}\triangleq\alpha_{n,k}c_{n,i}$. This results in the following compact, MoM-consistent expansion for the surface current:
$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum_{n=1}^{N_{b}}\sum_{i=1}^{M}\tilde{c}_{n,i,k}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right).$ (7)
This expression is structurally isomorphic to the MoM formulation in (2), with each $\tilde{c}_{n,i,k}$ acting as the learnable counterpart of each $I_{n,k}$ and $\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right)$’s serving as data-driven, adaptive generalizations of the fixed RWG basis functions. Critically, unlike MoM, which only solves for $I_{n,k}$’s against a fixed basis $\{{f}_{n}(\cdot)\}_{n=1}^{3}$, we regard KAN as a learnable generalization of the RWG basis that is deployed to preserve the MoM structure while enabling physics-adaptive basis functions.
III Proposed PhyKAN Framework
Figure 2: Reconstruction results for the surfaces of two objects with different variants of the proposed two-branch PhyKAN architecture.
In this section, we build upon the preceding formulation that recasts the MoM basis functions into a learnable KAN-based representation, and introduce PhyKAN, a physics-guided learning framework derived from the EFIE. The proposed artificial neural network, whose architecture is visualized in Fig. 1(a), has a dual-branch structure that enables adaptive basis learning by separately modeling local geometric features and global EM interactions.
III-A Neural Network Architecture
The local branch learns edge-wise basis representations by treating each RWG basis function as a node consistent with MoM, where an RWG basis is defined on the shared edge of two triangular faces. The input is a set of $M$ edge-nodes, each described by a geometric feature vector $\mathbf{g}_{m}\in\mathbb{R}^{d}$ encoding local geometry, where $d$ is the dimension of the feature vector that encodes local geometric information. As shown in Fig. 1(a), this descriptor is then mapped through a KAN block to produce a latent feature vector, as follows:
$\mathbf{h}^{(l)}_{m}\triangleq\mathrm{KAN}_{\boldsymbol{u}}\left(\mathbf{g}_{m}\right).$ (8)
where $\mathbf{h}^{(l)}_{m}\in\mathbb{R}^{D}$ denotes the $l$-th layer embedding. As shown in Fig. 1(b)), each KAN layer uses spline-based learnable activations (with optional $B$-spline or Fourier kernels) to enable adaptive basis shaping. The outputs of the three layers are fused to form the complete latent representation of each edge embedding $\mathbf{h}_{m}$, which serves as the learned local counterpart of the conventional RWG basis.
The local feature $\mathbf{H}_{k}$ of each $k$-th element-node is aggregated from its three associated edge local embeddings of the corresponding edges ($\mathbf{h}_{a}$, $\mathbf{h}_{b}$, and $\mathbf{h}_{c}$), as follows:
$\mathbf{H}_{k}=\frac{1}{3}\left(\mathbf{h}_{a}+\mathbf{h}_{b}+\mathbf{h}_{c}\right),$ (9)
providing a learnable counterpart of the RWG basis local expansion.
The global branch of PhyKAN models long-range electromagnetic interactions across the surface. In contrast to the local branch where each node corresponds to a mesh edge, here one element-node is defined per triangular element. The input to this branch is a set of $K$ element-nodes, where each node $k$ is represented by the centroid $\boldsymbol{p}_{k}\in\mathbb{R}^{3}$ of the $k$-th element, together with the incident direction $(\theta,\phi)$ of the scattered EM fields $\boldsymbol{E}_{\rm inc}(\boldsymbol{p}_{k})$’s. To incorporate EM priors into this branch, we initialize the weights of its first layer, $\mathbf{W}_{k,j}^{(0)}\in\mathbb{R}^{K\times K}$, using the dyadic Green’s function, as follows:
$\mathbf{W}_{k,j}^{(0)}=\mathbf{G}\left(\boldsymbol{p}_{k},\boldsymbol{p}_{j}\right),$ (10)
which encodes the physical interaction between element-nodes $k$ and $j$. To incorporate the dependence on incidence, the incident angles $(\theta,\phi)$ are mapped into a $D_{\rm ang}$-dimensional (with $D_{\rm ang}\triangleq 4L$) Fourier embedding, as follows:
$\boldsymbol{\gamma}(\theta,\phi)=\big[\sin(\omega_{1}\theta),\cos(\omega_{1}\theta),\ldots,\sin(\omega_{L}\phi),\cos(\omega_{L}\phi)\big],$ (11)
where $\{\omega_{\ell}\}_{\ell=1}^{L}$ are frequency parameters. In the sequel, the global coupling features and the latter $D_{\rm ang}$-directional embedding are concatenated and sequentially passed through two linear projections, yielding:
$\mathbf{F}_{k}=\mathbf{W}^{(2)}\left(\sigma\left(\mathbf{W}^{(1)}\left(\sum_{j=1}^{K}\mathbf{W}_{kj}^{(0)}\boldsymbol{p}_{j}\right)\oplus\boldsymbol{\gamma}(\theta,\phi)\right)\right)+\mathbf{b},$ (12)
where $\oplus$ denotes concatenation, $\sigma(\cdot)$ is an activation function, and $\mathbf{W}^{(1)}$, $\mathbf{W}^{(2)}$, and $\mathbf{b}$ are trainable parameters.
To finally predict surface currents, the local basis feature $\mathbf{H}_{k}$ and the global physical context $\mathbf{F}_{k}$ are combined as follows:
$\mathbf{z}_{k}=\mathbf{H}_{k}\oplus\mathbf{F}_{k},$ (13)
and are then projected to the $6$-dimensional output $\boldsymbol{J}_{\rm pre}(\boldsymbol{p}_{k})\in\mathbb{C}^{3\times 1}$ via a linear head.
III-B Physics-Informed Loss Function
Instead of directly minimizing the difference between predicted and reference currents, we adopt a residual-based loss function derived directly from the EFIE, which is defined as:
${L}=\frac{1}{N_{s}}\sum_{u=1}^{N_{s}}\left\|\sum_{j=1}^{K}\mathbf{G}(\boldsymbol{p}_{k},\boldsymbol{p}_{j})\boldsymbol{J}_{\rm pre}(\boldsymbol{p}_{k})-\boldsymbol{E}_{\rm inc}(\boldsymbol{p}_{k})\right\|_{2}^{2},$ (14)
where $N_{s}$ is the batch size and $\boldsymbol{p}_{k}$ is the center of the $k$-th triangular mesh element. This loss enforces the predicted surface currents to satisfy the EFIE, thereby ensuring physical consistency without requiring any supervised labels.
IV Numerical Results and Discussion
TABLE I: Our Proposed PhyKAN Framework’s Reconstruction Performance.
Dataset Cube Sphere Cone Assembly Body
$B$-spline PhyKAN 0.0098 0.0089 0.0084 0.0092
Fourier PhyKAN 0.0123 0.0072 0.0101 0.0104
Rational PhyKAN 0.0115 0.0107 0.0119 0.0145
KA-GCN PhyKAN 0.0061 0.0075 0.0060 0.0064
KA-GAT PhyKAN 0.0107 0.0119 0.0079 0.0088
To evaluate PhyKAN, MoM solutions are used as ground-truth for current reconstruction and RCS validation. The model inputs triangular mesh elements with EM descriptors and outputs induced currents. Tests are performed on four canonical PEC shapes [16] (cube, sphere, cone, assembly body) under a 1 GHz, unit-amplitude, vertical plane wave with $\theta,\phi\in[0^{\circ},180^{\circ}]$. Five variants assess the effect of KAN designs: three kernel activations ($B$-spline, Fourier, Rational) [18, 19, 20] and two graph KAN extensions (KA-GCN, KA-GAT) [21], used interchangeably in the local KAN blocks. Each is denoted [Kernel/Graph]-PhyKAN. DL-based CEM baselines either black-box the current mapping or only accelerate MoM without learnable bases; hence, we benchmark PhyKAN variants under a unified setting against MoM.
Figure 2 visualizes the surface reconstruction capability of five different variants of the proposed two-branch PhyKAN architecture. As observed, on the smooth geometry of the sphere, Fourier and KA-GCN PhyKAN accurately recover the global current distribution, while the $B$-spline and KA-GCN variants better capture sharp-edge behaviors on complex assemblies through localized and hierarchical aggregation. Quantitatively, Table I summarizes the mean-squared error (MSE) of each PhyKAN variant with respect to the MoM reference on the cube, sphere, cone, and assembly datasets. All variants show low reconstruction error against MoM, and the Fourier and KA-GCN configurations consistently achieve the highest accuracy across the canonical geometries. These results confirm that PhyKAN serves as a learnable, physics-adaptive basis, effectively overcoming the static rigidity of traditional RWG representations. This is actually attributed to its capability to tailor its current representation to the underlying geometry and field physics.
Figure 3: Bistatic RCS at $\phi=0^{\circ}$ for (a) the cube and (b) the cone objects, comparing different PhyKAN variants against the MoM reference.
TABLE II: Green’s Function Initialization (init.) Ablation for the Assembly Body Reconstruction.
Kernel MSE (no init.) MSE (Green init.)
KA-GCN PhyKAN 0.0123 0.0064
KA-GAT PhyKAN 0.0191 0.0088
To further assess the physical fidelity of the learned surface currents, the radar cross section (RCS) [24] is computed as a downstream validation in Fig. 3. It is shown that the predicted RCSs with all five considered PhyKAN variants agrees well with MoM references for both the cube and cone, even though this metric was never included as a training objective. This demonstrates that PhyKAN generalizes from near-field currents to physically consistent far-field responses, validating its physical consistency and confirming it avoids non-generalizable, black-box behavior.
In Table II, an ablation study, considering the assembly body, that directly validates the “Phy” component of the proposed PhyKAN framework is included. It can be seen that using Green’s function initialization significantly reduces reconstruction error. This confirms that embedding this physical prior is essential, resulting in strengthening both model accuracy and physical consistency. Overall, the results confirm that the proposed PhyKAN framework delivers geometry-adaptive and physically consistent surface current modeling, unifying data-driven flexibility with EM rigor.
V Conclusion
This letter introduced PhyKAN, a physics-guided EFIE-derived framework that enables a learnable basis for EM modeling. By generalizing the RWG basis into an adaptive and interpretable representation, the proposed PhyKAN framework unifies diverse kernel and structural extensions within a physically consistent architecture. Future work will focus on scaling the learnable basis to large-scale and broadband scenarios.

### v3

@@
-Learning the Basis: A Kolmogorov-Arnold Network Approach Embedding Green’s Function Priors
+Learning the Basis: A Kolmogorov-Arnold Network Approach Embedding Green’s Function Prior
 Abstract
-The Method of Moments (MoM) is constrained by the usage of static, geometry-defined basis functions, such as the Rao-Wilton-Glisson (RWG) basis.
-This letter reframes electromagnetic modeling around a learnable basis representation rather than solving for the coefficients over a fixed basis.
-We first show that the RWG basis is essentially a static and piecewise-linear realization of the Kolmogorov-Arnold representation theorem.
-Inspired by this insight, we propose PhyKAN, a physics-informed Kolmogorov-Arnold Network (KAN) that generalizes RWG into a learnable and adaptive basis family.
-Derived from the EFIE, PhyKAN integrates a local KAN branch with a global branch embedded with Green’s function priors to preserve physical consistency.
-It is demonstrated that, across canonical geometries, PhyKAN achieves sub-$10^{-2}$ reconstruction errors as well as accurate, unsupervised radar cross section predictions, offering an interpretable, physics-consistent bridge between classical solvers and modern neural network models for electromagnetic modeling.
+This letter proposes PhyKAN, a novel electromagnetics-informed learning framework that embeds the governing Electric Field Integral Equation (EFIE) directly into a dual-branch architecture for surface-current modeling.
+Inspired by the Kolmogorov-Arnold representation theorem, PhyKAN formulates a physics-informed framework where the local branch learns geometry-adaptive basis functions via Kolmogorov–Arnold Networks (KAN), while the global branch incorporates Green’s function priors to preserve long-range interactions and physical consistency.
+This design offers a unified formulation that generalizes classical Rao-Wilton-Glisson (RWG) basis into a data-driven, adaptive, and interpretable representation.
+It is demonstrated that, across canonical geometries, PhyKAN achieves reconstruction relative errors below 3%, preserves accurate RCS patterns, and remains robust across ablation, frequency, and mesh-resolution tests, with interpretable KAN-based complexity.
 I Introduction
@@
 This method expands the induced surface current using predefined basis functions, typically the Rao–Wilton–Glisson (RWG) basis [2].
-However, these basis functions are fixed by mesh geometry and are inherently static, a fact that prevents them from adapting to the spatial complexity of current distributions.
-This rigidity reveals a key limitation of MoM, as its basis functions are not learnable and cannot adapt to the underlying physics.
+Despite its success, MoM remains fundamentally non-learnable because both the mesh-prescribed basis space and the Green’s function operator kernel [22] are fixed, limiting its ability to adapt to geometric or physical complexity.
 To address the latter limitation, several directions have been explored.
@@
 On another front, instead of improving the basis itself, Adaptive Mesh Refinement (AMR) refines the discretization in regions of high error [8, 9].
-Although effective, these approaches remain largely hand-crafted, problem-specific, and fundamentally tied to predefined basis formulations.
+Although effective, these approaches remain hand-crafted and problem-specific, and they retain the fixed discretization structure of MoM, leaving the overall formulation fundamentally non-adaptive.
 Deep learning techniques are recently emerging as efficient tools in CEM problems, mainly along two directions: black-box current or field prediction without any explicit basis representation [10, 11], and MoM acceleration through surrogate models, operator learning, or iterative refinement, still relying on fixed basis functions [12, 13, 14].
@@
 KAN is flexible in both kernel choices [19, 20] and graph-based extensions [21] that embed geometric and physical interactions, making it a strong foundation for a physics-guided learnable basis for the MoM approach.
-In this letter, we reformulate MoM basis construction as a learnable function problem and show that KAN provides a theoretically grounded generalization of the RWG basis.
-In addition, a physics-guided EFIE-based framework, termed as PhyKAN, that embeds Green’s function [22] priors for physically consistent and adaptive current modeling, while remaining compact and interpretable, is introduced.
-The proposed learning framework accommodates multiple kernel functions and structural extensions, offering a unified bridge between classical integral equation solvers and neural network models.
+In this letter, we introduce PhyKAN, a dual-branch, physics-informed framework that embeds the EFIE directly into a learnable architecture.
+The key innovation is to replace the fixed Rao–Wilton–Glisson basis with KAN-based adaptive basis functions, allowing the surface-current representation to follow geometric and physical variations.
+Meanwhile, the global branch incorporates the Green’s function kernel as an explicit prior, lifting it into a higher-dimensional latent space to model long-range coupling while preserving EFIE consistency.
+By uniting KAN-driven basis adaptation with a Green’s-function-embedded global prior, PhyKAN offers a compact, physically grounded reformulation of EFIE–MoM within a modern neural architecture.
 II Surface Current as a Learning Problem
-Figure 1: (a) The architecture of the proposed PhyKAN comprising a local branch that encodes geometric features using a conventional KAN and global branch that incorporates Green’s function priors for physically consistent prediction of the induced surface current.
-(b) The structure of the conventional KAN [17], where learnable activation functions enable adaptive basis modeling.
-Let $\mathbf{G}\left(\cdot\right)\in\mathbb{C}^{3\times 3}$ denote the dyadic Green’s function in the three-dimensional free space, $\boldsymbol{J}(\cdot)\in\mathbb{C}^{3\times 1}$ the induced surface current, and $\boldsymbol{E}_{\rm inc}(\cdot)\in\mathbb{C}^{3\times 1}$ the incident field.
-By applying the boundary conditions of Perfectly Electrically Conducting (PEC) objects $\mathcal{D}$, the EFIE can be formulated $\forall\boldsymbol{p}\in\mathbb{R}^{3}$ as:
-$\int_{\boldsymbol{p}^{\prime}\in\mathcal{D}}\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)\boldsymbol{J}\left(\boldsymbol{p}^{\prime}\right)d\boldsymbol{p}^{\prime}=\boldsymbol{E}_{\rm inc}(\boldsymbol{p}).$ (1)
-The MoM discretizes this integral equation by expanding the unknown current $\boldsymbol{J}(\boldsymbol{p})$ using a set of predefined basis functions.
+Let $\mathbf{G}(\cdot)\in\mathbb{C}^{3\times 3}$ denote the dyadic Green’s function in three-dimensional free space, $\boldsymbol{J}(\cdot)\in\mathbb{C}^{3\times 1}$ the induced surface current, and $\boldsymbol{E}_{\rm inc}(\cdot)\in\mathbb{C}^{3\times 1}$ the incident field.
+Let $\boldsymbol{n}(\boldsymbol{p})$ denote the outward unit normal vector at $\boldsymbol{p}$ on the PEC surface $\mathcal{S}$.
+By applying the boundary conditions of a Perfectly Electrically Conducting (PEC) object with surface $\mathcal{S}$, the EFIE can be formulated $\forall\boldsymbol{p}\in\mathcal{S}$ as:
+$\mathbf{n}(\boldsymbol{p})\times\left[\mathbf{E}_{\text{inc }}(\boldsymbol{p})+\int_{\mathcal{S}}\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)\boldsymbol{J}\left(\boldsymbol{p}^{\prime}\right)\mathrm{d}\mathcal{S}^{\prime}\right]=\mathbf{0},\quad\boldsymbol{p}\in\mathcal{S}$ (1)
+Here, $\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)$ governs the electromagnetic interaction, while $\boldsymbol{J}\left(\boldsymbol{p}^{\prime}\right)$ is expanded through MoM basis functions.
+Together, these two elements define the core structure that PhyKAN reformulates into a learnable and physically consistent representation.
+II-A Learning the Basis
+The MoM discretizes this integral equation (1) by expanding the unknown current $\boldsymbol{J}(\boldsymbol{p})$ using a set of predefined basis functions.
 Typically, the RWG basis $\{{f}_{n}(\cdot)\}_{n=1}^{3}$ comprising fixed and geometry-defined functions [2] is used, yielding the following surface current approximation:
@@
 This geometry-fixed basis restricts the MoM’s adaptability to complex current distributions.
-Existing refinements, such as higher-order bases [3, 4, 5], the characteristic basis function method (CBFM) [7], and AMR[8, 9], partially alleviate this limitation but do not alter the static nature of the basis itself.
-Although recent deep learning works [10, 11, 12, 13, 14, 15, 16] attempt to either learn geometry-to-current mappings or accelerate MoM via surrogate models, they still operate on fixed basis functions, thus leaving the aforedescribed core limitation unresolved.
-This reveals a fundamental gap: a formulation that preserves the MoM structure while enabling adaptive basis functions is still missing.
-To this end, instead of solving for $I_{n,k}$ in (2) satisfying the EFIE in (1), we focus on learning an optimal, data-driven basis family $\left\{b_{n}(\cdot;\boldsymbol{u})\right\}$, yielding the following generalized expansion:
+In the classical MoM formulation, all degrees of freedom reside in the scalar coefficients $\{I_{n,k}\}$, whereas the basis shapes $\{f_{n}(\cdot)\}$ are entirely dictated by the mesh geometry and remain identical across different illuminations or target configurations.
+The RWG family is edge-based and divergence-conforming, which makes it physically sound but also rigid: improving expressiveness typically relies on mesh refinement or higher-order/macrobasis constructions, increasing the number of unknowns rather than enriching the basis itself.
+To introduce adaptability while preserving the MoM structure, we reinterpret basis construction itself as a learnable function approximation problem.
+Rather than only solving for fixed coefficients $I_{n,k}$ in (2) that scale a geometry-defined basis, we seek an optimal, data-driven basis family $\{b_{n}(\cdot;\boldsymbol{u})\}$ whose functional shape can adapt to the local geometry and electromagnetic response, while the expansion in (3) keeps the familiar MoM form, yielding the following generalized expansion:
 $\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum^{N_{b}}_{n=1}\alpha_{n,k}b_{n}(\boldsymbol{p}^{\prime}_{k};\boldsymbol{u}).$ (3)
@@
 Notably, setting $b_{n}=f_{n}$ and $\alpha_{n,k}=I_{n,k}$ recovers the classical MoM expansion in (2) as a special case.
-The latter problem finds its direct mathematical foundation in the Kolmogorov–Arnold representation theorem, which states that any multivariate function can be decomposed into finite sums of univariate functions [23]:
-$f\left(x_{1},\ldots,x_{d}\right)=\sum_{i=1}^{2d+1}\alpha_{i}\left(\sum_{j=1}^{d}\phi_{ij}\left(x_{j}\right)\right),$ (4)
-where $d$ is the input dimension, $\alpha_{i}$ are learnable weights, and $\phi_{ij}(\cdot)$’s are univariate nonlinear mappings.
-The KAN is the neural network architecture that operationalizes this theorem.
-To construct learnable basis functions, we embed the KAN parameterization into each $b_{n}(\cdot;\boldsymbol{u})$ as
-$b_{n}\left(\boldsymbol{p}^{\prime}_{k};\boldsymbol{u}_{n}\right)=\sum_{i=1}^{M}c_{n,i}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right),$ (5)
-where $\phi(\cdot)$ plays the same role as the univariate nonlinear mappings $\phi_{ij}(\cdot)$ in (4), and $\left\{\boldsymbol{w}_{n,i},\beta_{n,i},c_{n,i}\right\}$ are learnable parameters controlling each kernel’s orientation, offset, and amplitude, respectively.
-The term $\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right)$ performs a learnable linear projection of the spatial coordinate $\boldsymbol{p}^{\prime}_{k}$, allowing each kernel to adapt to different geometric directions and field variations.
+This formulation is enabled by the Kolmogorov–Arnold representation theorem [23], which expresses multivariate functions as compositions of univariate nonlinear mappings.
+Motivated by this structure, each basis function is parameterized using a KAN representation:
+$b_{n}\left(\boldsymbol{p}^{\prime}_{k};\boldsymbol{u}_{n}\right)=\sum_{i=1}^{M}c_{n,i}\phi\left(\boldsymbol{w}_{n,i}^{\top}\boldsymbol{p}_{k}^{\prime}+\beta_{n,i}\right),$ (4)
+where $\phi(\cdot)$ plays the same role as the univariate nonlinear mappings $\phi_{ij}(\cdot)$ in [23], and $\left\{\boldsymbol{w}_{n,i},\beta_{n,i},c_{n,i}\right\}$ are learnable parameters controlling each kernel’s orientation, offset, and amplitude, respectively.
+The term $\phi\left(\boldsymbol{w}_{n,i}^{\top}\boldsymbol{p}_{k}^{\prime}+\beta_{n,i}\right)$ performs a learnable linear projection of the spatial coordinate $\boldsymbol{p}^{\prime}_{k}$, allowing each kernel to adapt to different geometric directions and field variations.
 Substituting (4) into (3), yields the approximation:
-$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum_{n=1}^{N_{b}}\sum_{i=1}^{M}\alpha_{n,k}c_{n,i}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right).$ (6)
-To more closely mirror the MoM form in (2), we next absorb the coefficient into the spline weights via the definition $\tilde{c}_{n,i,k}\triangleq\alpha_{n,k}c_{n,i}$.
-This results in the following compact, MoM-consistent expansion for the surface current:
-$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum_{n=1}^{N_{b}}\sum_{i=1}^{M}\tilde{c}_{n,i,k}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right).$ (7)
-This expression is structurally isomorphic to the MoM formulation in (2), with each $\tilde{c}_{n,i,k}$ acting as the learnable counterpart of each $I_{n,k}$ and $\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right)$’s serving as data-driven, adaptive generalizations of the fixed RWG basis functions.
-Critically, unlike MoM, which only solves for $I_{n,k}$’s against a fixed basis $\{{f}_{n}(\cdot)\}_{n=1}^{3}$, we regard KAN as a learnable generalization of the RWG basis that is deployed to preserve the MoM structure while enabling physics-adaptive basis functions.
+$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum_{n=1}^{N_{b}}\sum_{i=1}^{M}\tilde{c}_{n,i,k}\phi\left(\boldsymbol{w}_{n,i}^{\top}\boldsymbol{p}_{k}^{\prime}+\beta_{n,i}\right),$ (5)
+with $\tilde{c}_{n,i,k}\triangleq\alpha_{n,k}c_{n,i}$ acting as the learnable counterpart of each $I_{n,k}$ and $\phi\left(\boldsymbol{w}_{n,i}^{\top}\boldsymbol{p}_{k}^{\prime}+\beta_{n,i}\right)$’s serving as data-driven, adaptive generalizations of the fixed RWG basis functions.
+The resulting expansion mirrors the MoM form in (2) while replacing the static RWG basis with a KAN-based, physics-adaptive representation.
+II-B Embedding Green’s Function Physics
+The dyadic Green’s function $\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)$ in (1) fully characterizes the electromagnetic interaction kernel in the EFIE and governs long-range coupling across the surface.
+In classical MoM, this operator is fixed and directly used to assemble the system matrix, leaving no room for learning enriched representations of global physics.
+To introduce learnability without violating the EFIE structure, PhyKAN incorporates $\mathbf{G}$ as an explicit physics prior and then lifts this operator into a higher-dimensional latent space.
+Formally, for the $k$-th element we construct
+$\mathbf{F}_{k}=\Phi\left(\mathbf{G}\left(\boldsymbol{p}_{k},\boldsymbol{p}_{k^{\prime}}\right)\right)$ (6)
+where $\Phi(\cdot)$ is a linear mapping initialized to satisfy $\Phi(\mathbf{G})=\mathbf{G}$, ensuring exact physical fidelity at the start of training.
+During optimization, $\Phi$ lifts the Green’s kernel $\mathbf{G}$ into a latent representation $\mathbf{F}_{k}\in\mathbb{C}^{D}$($D>3$) preserving the EFIE prior while enabling enriched, geometry-aware representations of long-range coupling.
+This high-dimensional lifting contrasts sharply with classical MoM, where the Green’s kernel is confined to a fixed $3\times 3$ dyadic structure.
+By embedding $\mathbf{G}$ as a prior and expanding its representational space, PhyKAN retains strict physical consistency while unlocking additional flexibility for capturing complex, nonlocal interactions.
 III Proposed PhyKAN Framework
-Figure 2: Reconstruction results for the surfaces of two objects with different variants of the proposed two-branch PhyKAN architecture.
-In this section, we build upon the preceding formulation that recasts the MoM basis functions into a learnable KAN-based representation, and introduce PhyKAN, a physics-guided learning framework derived from the EFIE.
+Figure 1: (a) PhyKAN consists of a KAN-based local branch that encodes geometry-adaptive basis features and a global branch that incorporates Green’s function priors for physically consistent modeling of long-range electromagnetic interactions.
+(b) The conventional KAN structure [17], where learnable activation functions form adaptive univariate kernels.
+In this section, we build upon the preceding formulation into PhyKAN, an EFIE-derived, physics-guided learning framework.
 The proposed artificial neural network, whose architecture is visualized in Fig.
@@
 III-A Neural Network Architecture
-The local branch learns edge-wise basis representations by treating each RWG basis function as a node consistent with MoM, where an RWG basis is defined on the shared edge of two triangular faces.
-The input is a set of $M$ edge-nodes, each described by a geometric feature vector $\mathbf{g}_{m}\in\mathbb{R}^{d}$ encoding local geometry, where $d$ is the dimension of the feature vector that encodes local geometric information.
-As shown in Fig.
-1(a), this descriptor is then mapped through a KAN block to produce a latent feature vector, as follows:
-$\mathbf{h}^{(l)}_{m}\triangleq\mathrm{KAN}_{\boldsymbol{u}}\left(\mathbf{g}_{m}\right).$ (8)
+The local branch learns edge-wise basis representations by treating each RWG basis function as a node, consistent with the MoM formulation in which each RWG function is defined on the shared edge of two triangular faces.
+The input is a set of $M$ edge-nodes, each described by a geometric feature vector $\mathbf{g}_{m}\in\mathbb{R}^{d}$ where the dimension $d$ of the feature vector includes edge length, adjacent-face normals, and relative edge orientation.
+As illustrated in Fig.
+1(a), the geometric descriptor is passed through a three-layer KAN encoder.
+Each KAN layer performs a nonlinear lifting of $\mathbf{g}_{m}$ through spline-based learnable activations:
+$\mathbf{h}^{(l)}_{m}\triangleq\mathrm{KAN}_{\boldsymbol{u}}\left(\mathbf{g}_{m}\right),$ (7)
 where $\mathbf{h}^{(l)}_{m}\in\mathbb{R}^{D}$ denotes the $l$-th layer embedding.
-As shown in Fig.
-1(b)), each KAN layer uses spline-based learnable activations (with optional $B$-spline or Fourier kernels) to enable adaptive basis shaping.
-The outputs of the three layers are fused to form the complete latent representation of each edge embedding $\mathbf{h}_{m}$, which serves as the learned local counterpart of the conventional RWG basis.
-The local feature $\mathbf{H}_{k}$ of each $k$-th element-node is aggregated from its three associated edge local embeddings of the corresponding edges ($\mathbf{h}_{a}$, $\mathbf{h}_{b}$, and $\mathbf{h}_{c}$), as follows:
+Each layer contains learnable linear projections and kernel activations (B-spline, rational, or Fourier), enabling the basis to adapt its local functional shape.
+To obtain a stable and expressive embedding, the three KAN outputs are fused via a residual linear aggregator:
+$\mathbf{h}_{m}=\sum_{l=1}^{3}{\eta}_{l}\odot\mathbf{h}_{m}^{(l)},$ (8)
+where ${\eta}_{l}$ are learnable per-layer weighting vectors that fuse multi-level KAN features into the final geometry-adaptive edge embedding.
+The local feature $\mathbf{H}_{k}$ of each $k$-th triangular element is aggregated from its three associated edge local embeddings of the corresponding edges ($\mathbf{h}_{a}$, $\mathbf{h}_{b}$, and $\mathbf{h}_{c}$), as follows:
 $\mathbf{H}_{k}=\frac{1}{3}\left(\mathbf{h}_{a}+\mathbf{h}_{b}+\mathbf{h}_{c}\right),$ (9)
@@
 The input to this branch is a set of $K$ element-nodes, where each node $k$ is represented by the centroid $\boldsymbol{p}_{k}\in\mathbb{R}^{3}$ of the $k$-th element, together with the incident direction $(\theta,\phi)$ of the scattered EM fields $\boldsymbol{E}_{\rm inc}(\boldsymbol{p}_{k})$’s.
-To incorporate EM priors into this branch, we initialize the weights of its first layer, $\mathbf{W}_{k,j}^{(0)}\in\mathbb{R}^{K\times K}$, using the dyadic Green’s function, as follows:
+To embed electromagnetic physics into this branch, the first-layer coupling weights between element-nodes are initialized directly with the dyadic Green’s function:
 $\mathbf{W}_{k,j}^{(0)}=\mathbf{G}\left(\boldsymbol{p}_{k},\boldsymbol{p}_{j}\right),$ (10)
-which encodes the physical interaction between element-nodes $k$ and $j$.
+which encodes the physical interaction between elements $k$ and $j$ as governed by the EFIE kernel.
+This initialization ensures that global message passing is grounded in the true nonlocal behavior of Maxwell’s equations.
 To incorporate the dependence on incidence, the incident angles $(\theta,\phi)$ are mapped into a $D_{\rm ang}$-dimensional (with $D_{\rm ang}\triangleq 4L$) Fourier embedding, as follows:
@@
 III-B Physics-Informed Loss Function
-Instead of directly minimizing the difference between predicted and reference currents, we adopt a residual-based loss function derived directly from the EFIE, which is defined as:
-${L}=\frac{1}{N_{s}}\sum_{u=1}^{N_{s}}\left\|\sum_{j=1}^{K}\mathbf{G}(\boldsymbol{p}_{k},\boldsymbol{p}_{j})\boldsymbol{J}_{\rm pre}(\boldsymbol{p}_{k})-\boldsymbol{E}_{\rm inc}(\boldsymbol{p}_{k})\right\|_{2}^{2},$ (14)
-where $N_{s}$ is the batch size and $\boldsymbol{p}_{k}$ is the center of the $k$-th triangular mesh element.
-This loss enforces the predicted surface currents to satisfy the EFIE, thereby ensuring physical consistency without requiring any supervised labels.
+For each incident field $\boldsymbol{E}^{(m)}_{\text{inc}}$, we compute the EFIE residual over all $K$ surface elements, rather than supervising the currents with MoM solutions.
+The residual for the $m$-th incident field is written as:
+$\mathcal{L}^{(m)}=\frac{1}{K}\sum_{k=1}^{K}\left\|\boldsymbol{n}(\boldsymbol{p}_{k})\times\left(\boldsymbol{E}^{(m)}_{\text{inc}}(\boldsymbol{p}_{k})+\sum_{j=1}^{K}\boldsymbol{G}(\boldsymbol{p}_{k},\boldsymbol{p}_{j})\boldsymbol{J}^{(m)}_{\text{pre}}(\boldsymbol{p}_{j})\right)\right\|_{2}^{2}$ (14)
+where $\boldsymbol{p}_{k}$, $\boldsymbol{p}_{j}$ denote the centers of the centers of the $k$-th and $j$-th mesh elements.
+This residual enforces the predicted current $\boldsymbol{J}^{(m)}_{\text{pre}}$ to satisfy the PEC boundary condition for the incident field $\boldsymbol{E}^{(m)}_{\text{inc}}$.
+The overall physics-informed loss is obtained by averaging over a batch of $N_{s}$ incident fields:
+$\mathcal{L}_{\text{EFIE}}=\frac{1}{N_{s}}\sum_{m=1}^{N_{s}}\mathcal{L}^{(m)}.$ (15)
+This formulation enables PhyKAN to be trained without MoM current labels, relying purely on the discretized EFIE for physical supervision.
 IV Numerical Results and Discussion
-TABLE I: Our Proposed PhyKAN Framework’s Reconstruction Performance.
-Dataset Cube Sphere Cone Assembly Body
-$B$-spline PhyKAN 0.0098 0.0089 0.0084 0.0092
-Fourier PhyKAN 0.0123 0.0072 0.0101 0.0104
-Rational PhyKAN 0.0115 0.0107 0.0119 0.0145
-KA-GCN PhyKAN 0.0061 0.0075 0.0060 0.0064
-KA-GAT PhyKAN 0.0107 0.0119 0.0079 0.0088
-To evaluate PhyKAN, MoM solutions are used as ground-truth for current reconstruction and RCS validation.
-The model inputs triangular mesh elements with EM descriptors and outputs induced currents.
-Tests are performed on four canonical PEC shapes [16] (cube, sphere, cone, assembly body) under a 1 GHz, unit-amplitude, vertical plane wave with $\theta,\phi\in[0^{\circ},180^{\circ}]$.
-Five variants assess the effect of KAN designs: three kernel activations ($B$-spline, Fourier, Rational) [18, 19, 20] and two graph KAN extensions (KA-GCN, KA-GAT) [21], used interchangeably in the local KAN blocks.
-Each is denoted [Kernel/Graph]-PhyKAN.
+To evaluate PhyKAN, we assess its effectiveness, physical consistency, and adaptive behavior as a learnable basis.
+High-fidelity MoM solutions on a fine mesh ($\lambda/10$), the standard CEM reference [1], serve as ground truth for both surface-current reconstruction and Radar Cross Section (RCS) validation.
+The model takes triangular mesh elements with EM descriptors and predicts induced currents.
+Experiments cover four canonical PEC targets [16] (cube, sphere, cone, assembly body) and a complex SLICY object [25] under a 1 GHz, unit-amplitude, vertically polarized plane wave with $\theta,\phi\in[0^{\circ},180^{\circ}]$.
 DL-based CEM baselines either black-box the current mapping or only accelerate MoM without learnable bases; hence, we benchmark PhyKAN variants under a unified setting against MoM.
-Figure 2 visualizes the surface reconstruction capability of five different variants of the proposed two-branch PhyKAN architecture.
-As observed, on the smooth geometry of the sphere, Fourier and KA-GCN PhyKAN accurately recover the global current distribution, while the $B$-spline and KA-GCN variants better capture sharp-edge behaviors on complex assemblies through localized and hierarchical aggregation.
-Quantitatively, Table I summarizes the mean-squared error (MSE) of each PhyKAN variant with respect to the MoM reference on the cube, sphere, cone, and assembly datasets.
-All variants show low reconstruction error against MoM, and the Fourier and KA-GCN configurations consistently achieve the highest accuracy across the canonical geometries.
-These results confirm that PhyKAN serves as a learnable, physics-adaptive basis, effectively overcoming the static rigidity of traditional RWG representations.
-This is actually attributed to its capability to tailor its current representation to the underlying geometry and field physics.
-Figure 3: Bistatic RCS at $\phi=0^{\circ}$ for (a) the cube and (b) the cone objects, comparing different PhyKAN variants against the MoM reference.
-TABLE II: Green’s Function Initialization (init.) Ablation for the Assembly Body Reconstruction.
-Kernel MSE (no init.) MSE (Green init.)
-KA-GCN PhyKAN 0.0123 0.0064
-KA-GAT PhyKAN 0.0191 0.0088
-To further assess the physical fidelity of the learned surface currents, the radar cross section (RCS) [24] is computed as a downstream validation in Fig.
-3.
-It is shown that the predicted RCSs with all five considered PhyKAN variants agrees well with MoM references for both the cube and cone, even though this metric was never included as a training objective.
-This demonstrates that PhyKAN generalizes from near-field currents to physically consistent far-field responses, validating its physical consistency and confirming it avoids non-generalizable, black-box behavior.
-In Table II, an ablation study, considering the assembly body, that directly validates the “Phy” component of the proposed PhyKAN framework is included.
-It can be seen that using Green’s function initialization significantly reduces reconstruction error.
-This confirms that embedding this physical prior is essential, resulting in strengthening both model accuracy and physical consistency.
-Overall, the results confirm that the proposed PhyKAN framework delivers geometry-adaptive and physically consistent surface current modeling, unifying data-driven flexibility with EM rigor.
+Figure 2: Surface current reconstruction on (a) the sphere and (b) the assembly body.
+Columns show the MoM reference, PhyKAN prediction, absolute error, KAN complexity maps, and the corresponding RCS comparison.
+TABLE I: Relative MSE ($10^{-2}$) of PhyKAN and ablated variants.
+Method Cube Sphere Cone Assembly Body
+PhyKAN (no physics init) 3.42 2.95 3.88 6.73
+PhyKAN (MLP Basis) 4.75 3.89 5.26 8.94
+PhyKAN 1.21 0.94 1.28 1.84
+TABLE II: Relative MSE ($10^{-2}$) of PhyKAN at different frequencies.
+Frequency Cube Sphere Cone Assembly
+1 GHz 1.21 0.94 1.28 1.84
+2 GHz 1.46 1.07 1.53 2.21
+5 GHz 1.89 1.33 1.98 2.67
+Figure 2 shows the reconstruction results.
+Across both the sphere and the assembly body, PhyKAN closely matches the MoM current distributions, with absolute errors remaining small and localized.
+The KAN complexity maps quantify the average activation magnitude of all kernels in KAN blocks across the three edges of each element, yielding higher values near geometric discontinuities and thus highlighting PhyKAN’s adaptive allocation of $b_{n}(\cdot;\boldsymbol{u})$ compared with the fixed RWG basis.
+The far-field RCS results indicate that PhyKAN matches MoM patterns despite receiving no RCS supervision.
+Table I reports the relative MSE for the ablation study.
+Removing the Green’s function initialization leads to higher errors across all geometries, while replacing the KAN basis with a plain MLP causes an even larger degradation.
+The full PhyKAN model achieves the best accuracy, demonstrating the complementary benefits of physics-based initialization and the learnable KAN basis.
+Table II summarizes PhyKAN’s accuracy when trained at 1, 2, and 5 GHz.
+The error increases only slightly with frequency due to finer high-frequency current oscillations, and remains stable across all targets, with a modest 1.3–1.6× rise from 1 to 5 GHz.
+Fig.
+3 evaluates performance under various mesh resolutions, using relative MSE with $\lambda/25$ MoM result as reference.
+In coarse-mesh regimes (e.g., $\lambda$, $\lambda/2.5$, $\lambda/5$), PhyKAN substantially reduces MSE on coarse meshes by compensating for geometric undersampling with its learnable basis.
+MoM overtakes only at very fine resolutions where computational cost dominates, making PhyKAN most valuable in coarse-mesh regimes.
+Figure 3: Relative MSE ($10^{-2}$) of PhyKAN and MoM under different mesh resolutions, computed with respect to the $\lambda$/25 MoM reference.
 V Conclusion
-This letter introduced PhyKAN, a physics-guided EFIE-derived framework that enables a learnable basis for EM modeling.
-By generalizing the RWG basis into an adaptive and interpretable representation, the proposed PhyKAN framework unifies diverse kernel and structural extensions within a physically consistent architecture.
+This letter introduced PhyKAN, an EFIE-derived framework that replaces the fixed RWG basis with an adaptive learnable representation while embedding the Green’s function as a physics prior, forming a unified and physically consistent architecture for electromagnetic modeling.
 Future work will focus on scaling the learnable basis to large-scale and broadband scenarios.